Use synthetic division to find P(–2) for P(x) = x4 + 9x3 - 9x + 2 .
A. –2
B. 0
C. –36
D. 68

Answer by ewatrrr(24785)

 
Hi,
P(x) = x4 + 9x3 - 9x + 2
P(-2) = 16 -72 +18 + 2 = -36